A birds flies 3.0m west and then flies 5.0m south.


a) Draw the bird resultant displacement (d).


b) |d|


c) Direction of the bird’s resultant


1
Expert's answer
2021-11-02T10:01:31-0400

a)



b)

"|d|=\sqrt{3^2+5^5}=\sqrt{34}" m


c)

"\alpha=arctan(5/3)=59\degree" south of west